18:19 Start or run 7: In beam has been optimzed up to now by Zoran and Claude. ~7E7 particles at injection. From now on, only the injection beamline is optimized by Operators.
Detectors IN (left arm pot =18.86 mm, right arm pot = 19.80 mm)
No Target
Beam is in the ring for ~6 seconds and total cycle time between injections is ~8 seconds
Injection beamline is still in process of optimization by operators
21:30
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
23:30
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
01:30
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
Beam and Electron Coller failure (~00:35); Beam back (~01:08); Cooler back (~01:20)
03:00
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
Beam and Electron Coller failure (~01:35); Beam back (~01:55); Cooler back (~01:55)
04:40
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
06:00
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
07:00
Merger OK, Tape server OK
ADC data item statistics looks as before. OK
FEE temperatures look OK
System checks OK
Leakage current stable
|